# Zero-Shot Hyperspectral Pansharpening Using Hysteresis-Based Tuning for Spectral Quality Control

### v1

@@ IV-D Full-resolution quality assessment
$D_{\lambda}$ is a solid indicator of quality, as it is computed against a well defined and unambiguous reference. Measuring spatial quality is much more challenging, which is why we consider two alternative indices. In fact, neither of them is without flaws. For example, $D_{S}$ quantifies the distance between the PAN image and the optimal linear combination of the pansharpened spectral bands that best approximates it [92]. This is physically motivated by the fact that a high-resolution hyperspectral (HS) sensor would capture narrowband signals that collectively contribute to the broadband response of the PAN sensor. Therefore, an ideal set of sharpened HS bands should be able to reconstruct the PAN image through an appropriate linear combination. However, if one band (among hundreds) is an exact copy of the PAN, then $D_{S}=0$, no matter what the other bands look like. This happens for several entries in Tab. VIII which, far from signaling perfect pansharpening, correspond typically to degenerate solutions. While such cases are rarely seen with MS images, they occur more easily with HS images. Similar warnings apply to the second indicator. In fact, $D_{\rho}$ is defined as the average of the local (calculated on a $\sigma{\times}\sigma$ domain) correlation coefficient between PAN and the pansharpened HS bands [85]. From a physical perspective, this metric evaluates the extent to which a $\sigma{\times}\sigma$ patch from any band of $\widehat{\mathbf{\mathbf{H}}}$ can be linearly predicted from the corresponding patch in the PAN image. Consequently, $D_{\rho}$ directly reflects the degree of spatial alignment between the pansharpened product $\widehat{\mathbf{\mathbf{H}}}$ and the PAN image $P$ at fine spatial scales. Therefore, it vanishes only when there is perfect correlation between all bands and the PAN. However, this is not desirable for bands that have no spectral overlap with the PAN and may differ significantly from it. Although both indices have their limitations, we believe that their combined use offers a more comprehensive assessment of spatial consistency. Keeping this in mind, we comment on the numerical results with some caution and await confirmation from the visual analysis to draw firm conclusions. In particular, the proposed method performs quite well in terms of spatial quality indices, especially when considering $D_{\rho}$. Together with the excellent results of $D_{\lambda}$, this speaks of a reliable and well-balanced method. This does not apply to TV, however, whose performance drops dramatically, suggesting questionable spatial quality. On the contrary, MF is characterized by excellent spatial quality indices but a very bad $D_{\lambda}$. Both these methods, and many others, seem unable to achieve a good balance between spectral and spatial fidelity. Overall, the proposed method compares quite favorably with all DL-based references. Only R-PNN provides a better $D_{\rho}$ on average, but this is also due to the perfect alignment between the metric and the spatial loss of R-PNN, a property that does not hold with the new spatial loss used in the proposal. The numerical results for the Sentinel-2/PRISMA dataset are reported in Table IX. In this experiment, it is evident that none of the implemented methods, except for the proposed one, achieves a top-5 ranking in both spectral quality ($D_{\lambda}$) and spatial quality ($D_{S}$ and/or $D_{\rho}$).
@@ IV-E Visual results
In Fig. 10 and Fig. 11 we show results for two closeups from the FR Cagliari and the FR Tabasco datasets. Following [42], we show two different spectral triplets of bands for false-RGB composition, from the visible and NIR-SWIR spectral ranges. Given the large number of reference methods, we limit attention to the top-5 $D_{S}$ and top-5 $D_{\rho}$ methods (based on averages from Tab. VIII) plus TV, top-1 $D_{\lambda}$, for a total of 10 methods including the proposed one. [^3] The full set of results is available on the web repository at https://github.com/giu-guarino/rho-PNN.. Figure 12 presents the results for the Sentinel-2/PRISMA dataset. As in Figures 10 and 11, we include 10 methods following the same selection criteria, with the exception of replacing TV with PCA-Z-PNN (the second-best performer in terms of $D_{\lambda}$ after the proposed method). Together with the images, we also report the spectral quality ($D_{\lambda}$) and spatial quality ($D_{S}/D_{\rho}$) scores computed over all bands of the full scene. Additionally, Figure 13 presents the visual results of the top-5 methods in terms of $Q^{2n}$ on the RR Udine dataset, along with their corresponding error maps. As in previous cases, we also report the evaluation metrics ($ERGAS$, $SAM$, and $Q^{2n}$) computed over all spectral bands and the full scene.
@@
Figure 10: Pansharpening results on the FR Cagliari dataset (240$\times$360 close-up): PAN image (top left) with paired lower resolution HS input (bands: 663, 560 and 466 nm) followed by a selection of pansharpening results. The input HS and all results are also displayed, on even rows, using a selection of bands from the NIR-SWIR range (1943, 1261 and 832 nm). For each result the corresponding scores “$D_{\lambda}$, $D_{S}/D_{\rho}$” are reported (from Tab. VIII).
@@
Figure 11: Pansharpening results on the FR Tabasco dataset (240$\times$360 close-up): PAN image (top left) with paired lower resolution HS input (bands: 663, 560 and 466 nm) followed by a selection of pansharpening results. The input HS and all results are also displayed, on even rows, using a selection of bands from the NIR-SWIR range (1943, 1261 and 832 nm). For each result the corresponding scores “$D_{\lambda}$, $D_{S}/D_{\rho}$” are reported (from Tab. VIII).
@@
Figure 12: Pansharpening results on the Sentinel-2 / PRISMA dataset (240$\times$360 close-up): Simulated PAN image obtained from Sentinel-2 RGB bands (top left) with paired lower resolution PRISMA HS input (bands: 663, 560 and 466 nm) followed by a selection of pansharpening results. The input HS and all results are also displayed, on even rows, using a selection of bands from the NIR-SWIR range (1943, 1261 and 832 nm). For each result the corresponding scores “$D_{\lambda}$, $D_{S}/D_{\rho}$” are reported (from Tab. VIII).
GT HyperPNN HSpeNet DIP-HyperKite Hyper-DSNet $\rho$-PNN
[…]
Figure 13: Pansharpening results on the RR Udine dataset (120$\times$180 close-up) and error maps (Top 5 $Q2^{n}$). RGB (a) and False color (b) results are displayed. For each result the corresponding scores “ERGAS, SAM, $Q2^{n}$” are reported.
  • Results with low $D_{\rho}$ are generally sharper, resembling closely the spatial structures of the PAN. This is true for both PRISMA FR scenes and Sentinel-2/PRISMA scene and in both the visible and NIR-SWIR spectral ranges. In particular, this is the case of BT-H, MF, MTF-GLP-HPM, R-PNN and proposed.
  • Results with low $D_{S}$ reproduce the PAN spatial structures faithfully in the visible spectrum, but seem to fail outside that range. This is the case of BDSD-PC, HySURE and DIP-HyperKite form PRISMA FR scene, whereas GSA, BDSD-PC, and PRACS are the cases for Sentinel-2/PRISMA scene. BT-H seems an exception, characterized also by a low $D_{\rho}$, however it presents a large spectral distortion.
  • For PRISMA FR scenes, the top-$D_{\lambda}$ method, TV, whose spectral quality level is nearly the same as the proposed one, does not provide effective sharpening, providing blurred images.
For the RR evaluation framework, we selected image crops from the Udine area (see Fig. 13), displaying results for the visible bands in the top row and the NIR–SWIR bands in the bottom row. The left image shows the ground truth (i.e., the original HS image), followed by the pansharpening results of the five best-performing methods according to the Q2n metric, and finally, the output of our proposed method. The visual quality of the results produced by the top-5 algorithms is generally comparable with that of our method, confirming the trends observed in the quantitative results reported in Tab. VII. However, a closer inspection reveals that our method tends to preserve spectral fidelity more effectively. In contrast, the output of HyperPNN exhibits noticeable color artifacts. Although some of the other methods provide slightly sharper spatial details, they also introduce visible spectral aberrations, particularly in homogeneous regions such as the sea.
Based on the above observations, we single out three methods, MTF-GLP-HPM, R-PNN and our proposal, that appear to be both robust and effective, performing uniformly well on all scenes and all displayed spectral bands in terms of both spectral and spatial quality. Samples of these methods are shown close to each other but the differences are so small that further examination seems like a futile exercise that could lead to claims related only to the choice of scene/bands/clips. On the other hand, all three methods have very similar spatial quality indicators. However, we point out that the proposed method has a much smaller average spectral distortion (0.0033, in Tab.VIII and 0.0013 in Tab.IX) than these competitors (0.0098 and 0.0090, respectively, in Tab.VIII, and 0.0039 and 0.0032, respectively in Tab.IX). Furthermore, only the proposed method ensures uniform quality across all bands, a feature not captured by the global indicators.
In summary, the proposed method demonstrates more consistent and robust performance compared to current SotA. It effectively preserves spectral fidelity without compromising spatial quality, and unlike other deep learning-based approaches, it does not require any training data for pansharpening inference. Furthermore, as confirmed by an additional test involving a PAN image acquired from a different sensor on a different date, the method exhibits a strong degree of generalizability. On the downside, under the current hyperparameter configuration, the approach entails a relatively high computational cost compared to both model-based and several deep learning-based techniques. However, by relaxing the spectral quality constraints (via $\gamma_{\text{high}}$ and $\gamma_{\text{low}}$), it is possible to accelerate the algorithm significantly.
@@ V Conclusions
Advantages
  • It works from scratch (Only PCA-Z-PNN does the same);
  • It works with an arbitrary number of bands (Only PCA-Z-PNN and R-PNN do the same);
  • It ensures a good generalization ability across diverse sensors;
  • It works at full resolution: No downgrade nor simulation operations are needed (same as R-PNN and PCA-Z-PNN);
  • It ensrures a controllable and uniform level of spectral control along all wavelengths.
Drawbacks
  • It is slower than pre-trained methods at inference time;

### v2

@@ IV-D Full-resolution quality assessment
$D_{\lambda}$ is a solid indicator of quality, as it is computed against a well defined and unambiguous reference. Measuring spatial quality is much more challenging, which is why we consider two alternative indices. In fact, neither of them is without flaws. For example, $D_{S}$ measures the distance between the PAN and the best (closest to the PAN) linear combination of the pansharpened spectral bands [90]. Therefore, if one band (among hundreds) is an exact copy of the PAN, then $D_{S}=0$, no matter what the other bands look like. This happens for several entries in Tab. VIII which, far from signaling perfect pansharpening, correspond typically to degenerate solutions. While such cases are rarely seen with MS images, they occur more easily with HS images. Similar warnings apply to the second indicator. In fact, $D_{\rho}$ is defined as the average of the local correlation coefficient between PAN and the pansharpened HS bands [83]. From a physical perspective, this metric evaluates the extent to which a $\sigma{\times}\sigma$ patch from any band of $\widehat{\mathbf{\mathbf{H}}}$ can be linearly predicted from the corresponding patch in the PAN image. Consequently, $D_{\rho}$ directly reflects the degree of spatial alignment between the pansharpened product $\widehat{\mathbf{\mathbf{H}}}$ and the PAN image $P$ at fine spatial scales. Therefore, it vanishes only when there is perfect correlation between all bands and the PAN. However, this is not desirable for bands that have no spectral overlap with the PAN and may differ significantly from it.
Keeping this in mind, we comment on the numerical results with some caution and await confirmation from the visual analysis to draw firm conclusions. In particular, the proposed method performs quite well in terms of spatial quality indices, especially when considering $D_{\rho}$. Together with the excellent results of $D_{\lambda}$, this speaks of a reliable and well-balanced method. This does not apply to TV, however, whose performance drops dramatically, suggesting questionable spatial quality. On the contrary, MF is characterized by excellent spatial quality indices but a very bad $D_{\lambda}$. Both these methods, and many others, seem unable to achieve a good balance between spectral and spatial fidelity. Overall, the proposed method compares quite favorably with all DL-based references. Only R-PNN provides a better $D_{\rho}$ on average, but this is also due to the perfect alignment between the metric and the spatial loss of R-PNN, a property that does not hold with the new spatial loss used in the proposal.
@@ IV-E Visual results
In Fig. 10 and Fig. 11 we show results for two closeups from the FR Cagliari and the FR Tabasco datasets. Following [40], we show two different spectral triplets of bands for false-RGB composition, from the visible and NIR-SWIR spectral ranges. Given the large number of reference methods, we limit attention to the top-5 $D_{S}$ and top-5 $D_{\rho}$ methods (based on averages from Tab. VIII) plus TV, top-1 $D_{\lambda}$, for a total of 10 methods including the proposed one. [^3] The full set of results is available on the web repository at
https://github.com/giu-guarino/rho-PNN.
Together with the images, we also report the spectral ($D_{\lambda}$) and spatial ($D_{S}$ and $D_{\rho}$) quality scores computed over all bands of the full scene.
@@
Figure 10: Pansharpening results on the FR Cagliari dataset (240$\times$360 close-up): PAN image (top left) with paired lower resolution HS input (bands: 663, 560 and 466 nm) followed by a selection of pansharpening results. The input HS and all results are also displayed, on even rows, using a selection of bands from the NIR-SWIR range (1943, 1261 and 832 nm). For each result the corresponding scores “($D_{\lambda}$), ($D_{S}$, $D_{\rho}$)” are reported (from Tab. VIII).
@@
Figure 11: Pansharpening results on the FR Tabasco dataset (240$\times$360 close-up): PAN image (top left) with paired lower resolution HS input (bands: 663, 560 and 466 nm) followed by a selection of pansharpening results. The input HS and all results are also displayed, on even rows, using a selection of bands from the NIR-SWIR range (1943, 1261 and 832 nm). For each result the corresponding scores “($D_{\lambda}$), ($D_{S}$, $D_{\rho}$)” are reported (from Tab. VIII).
GT HyperPNN HSpeNet DIP-HyperKite Hyper-DSNet $\rho$-PNN
[…]
Figure 12: Pansharpening results on the RR Udine dataset (120$\times$180 close-up) for visible bands (a) and NIR-SWIR bands (b). In both cases, ground truth and pansharpening results (top row) with error maps (bottom row). For each result the corresponding scores “(ERGAS, SAM, $Q2^{n}$)” are reported.
  • Results with low $D_{\rho}$ are generally sharper, resembling closely the spatial structures of the PAN. This is true for scenes and in both the visible and NIR-SWIR spectral ranges. In particular, this is the case of BT-H, MF, MTF-GLP-HPM, R-PNN and proposed.
  • Results with low $D_{S}$ reproduce the PAN spatial structures faithfully in the visible spectrum, but seem to fail outside that range. This is the case of BDSD-PC, HySURE and DIP-HyperKite. BT-H seems an exception, characterized also by a low $D_{\rho}$, however it presents a large spectral distortion.
  • The top-$D_{\lambda}$ method, TV, whose spectral quality level is nearly the same as the proposed one, does not provide effective sharpening, providing blurred images.
Based on the above observations, we single out three methods, MTF-GLP-HPM, R-PNN and our proposal, that appear to be both robust and effective, performing uniformly well on both scenes and all displayed spectral bands in terms of both spectral and spatial quality. Samples of these methods are shown close to each other but the differences are so small that further examination seems like a futile exercise that could lead to claims related only to the choice of scene/bands/clips. On the other hand, all three methods have very similar spatial quality indicators. However, we point out that the proposed method has a much smaller average spectral distortion (0.0033, see Tab.VIII) than these competitors (0.0098 and 0.0090, respectively). Furthermore, only the proposed method ensures uniform quality across all bands, a feature not captured by the global indicators.
We complete this analysis by showing, in Fig. 12, some error maps obtained by subtracting the ground truth (GT) from the pansharpened image. In this case, we use the Udine dataset, where some interesting phenomena are visible, and work at reduced resolution, where the original HS component can be used as GT. To save space only 5 methods are considered, the proposed one and the best four competitors according to the Q2n metric. For bands in the visible range (top) only small errors occur, and the error maps appear almost uniformly gray, with the exception of HyperPNN where some color aberrations are visible. As expected, larger errors appear for the NIR–SWIR bands (bottom) that do not overlap the PAN. Interestingly, some sharp structures visible in the GT (probably, waterways) are almost completely lost in the fused images, uniformly for all methods. On the other hand, such structures are hardly distinguishable in the visible bands and therefore in the PAN, suggesting that the relevant information may be truly missing in this case.
IV-F Generalization Ability
As already said, $\rho$-PNN does not require any preliminary training phase but adapts on the fly to the target data. Therefore, we expect it to generalize smoothly to new sources, a key requirement of data-driven methods. To investigate this issue experimentally we would need an alternative dataset, but we could find none available online, so we built a new one ad hoc, taking advantage of the good coverage and wide availability of Copernicus Sentinel-2 imagery. We found a Sentinel-2 image covering the same area as our PRISMA Kansas dataset with a temporal shift of one day. We used the RGB (665, 560 and 490 nm) high-resolution bands of the Sentinel-2 image to synthesize a 10-meter resolution PAN, and co-registered it with the PRISMA HS dataset. Finally, to preserve the resolution ratio of 6 used in the toolbox [42] for all methods, the hyperspectral data were downscaled by a factor of 2, resulting in a spatial resolution of 60 meters.
On this heterogeneous dataset we repeated the same experiments carried out on the homogeneous PRISMA datasets working, however, only at full resolution, for lack of sufficient data. The Results are reported in Table IX and Figure 13 with the by now usual format, where comparison methods are selected based on the results of Table IX. Both numerical and visual results confirm our expectations: $\rho$-PNN keeps working very well and is the only method that achieves top-5 ranking for both spectral and spatial quality. On the contrary, several deep learning-based methods appear to struggle with these new data, especially in terms of spectral quality. This is especially obvious for HyperPNN, with pansharpened images that are severely corrupted, see Figure 13. Model-based methods, on the other hand, are not significantly affected by the data mismatch. In extreme synthesis, these results show that the proposed method keeps ensuring a very good and stable performance also on this new dataset, confirming the expected good generalization ability.
@@ IV-F Generalization Ability
Figure 13: Pansharpening results on the synthetic Sentinel-2/PRISMA dataset (240$\times$360 close-up): Simulated PAN image obtained from Sentinel-2 RGB bands (top left) with paired reduced resolution PRISMA HS input (bands: 663, 560 and 466 nm) followed by a selection of pansharpening results. The input HS and all results are also displayed, on even rows, using a selection of bands from the NIR-SWIR range (1943, 1261 and 832 nm). For each result the corresponding scores “($D_{\lambda}$), ($D_{S}$, $D_{\rho}$)” are reported (from Tab. IX).
IV-G Discussion
This large body of experimental results fully confirms our expectations about the proposed method. For the sake of clarity, we summarize here its pros and cons, also with reference to competing methods.
From a structural point of view, $\rho$-PNN presents several qualifying points, some of which are shared only with its ancestor R-PNN, and with PCA-Z-PNN:
  • it runs from scratch, without the need for prior training: this is obtained by means of a longer tuning run on the first band that affects only marginally the overall computational cost due to the very large number of bands;
  • it works with an arbitrary number of bands: as the algorithm works band-wise with no joint processing of the HS stack, the number of spectral bands does not need to be fixed to build the model;
  • it operates at full resolution with no resolution downgrading or simulation operations: as the proposed loss is fully unsupervised and based of input-output self-consistency, traditional resolution downgrade protocols (e.g., [82]) to give rise to labeled, but synthetic, dataset are avoided.
Furthermore, in terms of performance
  • it is on par with the best SotA methods, both model-based and data-driven;
  • it guarantees a uniformly high spectral quality at all wavelengths, as only the TV method does;
  • it shows excellent generalization ability.
On the down side, $\rho$-PNN has a relatively high computational cost compared to model-based techniques and even several data-driven techniques. This cost can be significantly reduced, however, by slightly relaxing the spectral quality constraints.